# Doomed Worlds I: No new evidence for orbital decay in a long-term survey of 43 ultra-hot Jupiters

### v1

@@ 5.2.1 Some examples of composite times in this work
In Section 6.4 we discuss WASP-121 b, where a single composite point, in this case an ephemeris time, is primarily responsible for the non-zero value. In that case, we do not have access to either the individual times or the associated light curves that were used to calculate , making its impact hard to evaluate.
CoRoT-2 b, meanwhile, contains an example of a stacked midtime. The time series from the CoRoT satellite mission encompasses 82 consecutive, high-quality individual transits, but the data were first published as a single stacked midtime assigned to the midtime of the first transit, with [BJD] (Alonso et al., 2008). In this work we reanalyzed the original photometry so that we could individually fit each of the 82 transits (see Table 5), which had an average error of -0.00020, +0.00023 d in the timing of each transit (50% larger than the error assigned to the single stacked midtime). In an experiment, we found significant shifts in the preferred ephemeris model using the stacked midtime rather than the 82 individual midtimes: fell from 33 to -2 and went from 2- positive curvature to negative curvature that is indistinguishable from zero. In this particular case, since each of the individual transits could be well fit to a light curve model on its own, there is no need to use a stacked transit midtime.
We also found that the CoRoT-2 b mission data were being double-counted in Ivshina & Winn (2022). Both the stacked composite time (derived from all 82 transits) from Alonso et al. (2008) and 79 of the 82 individual transit light curves are included in the table in Ivshina & Winn (2022). This duplication had two causes: (1) The individual transit midtimes came from Öztürk & Erdem (2019), which did not analyze the first three transits taken at the longer 512 s cadence; but since the stacked transit was assigned to the midtime of the first (omitted) transit, it appeared to be at a unique epoch. (2) A typo in Ivshina & Winn (2022) caused the omission of the “” tag to indicate that the Alonso et al. (2008) time was derived from more than one transit. In this case, double-counting the CoRoT data had only a modest downward effect on (decreasing it by 6).
@@ 5.3 Some words of caution on timing systems
Eastman et al. (2010) provides a thorough discussion of the various timing systems in use in modern astronomy. Briefly, to remove the effects of both the motion of the Earth around the Sun and the motion of the Sun around the solar system barycenter, a dynamical time system must be used to avoid introducing cyclical variations in an astronomical time series. The recommended standard for reporting exoplanet transit times is Barycentric Julian Dates (BJD) using the Barycentric Dynamical Time (TDB) scale. Critically, for precise timing analyses, all times must be converted into the same timing system before analysis so as to avoid introducing systematic errors.
@@
The difference between the solar system barycenter and the heliocenter is small (no more than a few seconds, Eastman et al., 2010) and BJD and HJD have sometimes been used interchangeably, since almost no transit midtimes are known to better than a few seconds. This practice would reasonable if the times were also both in the same dynamical system – but BJD times are almost always given in the TDB system, while HJD times are usually (though not always) in the UTC system. The difference between UTC and TDB is currently 69.184 s (in 2024). The UTC-TDB offset changes with the irregular addition of leap seconds that depends on the precise rotation of the Earth, and the last leap second was added on 2016 December 31555https://www.nist.gov/pml/time-and-frequency-division/time-realization/leap-seconds. Thus it is critical to report the full time system used, and herein lies some difficulties with literature light curves. Often, times are reported as BJD only, with BJD/TDB implied, though occasionally light curves will be reported in BJD/UTC. The table of transits from Ivshina & Winn (2022) reports the timing system used for each light curve, usually either BJD/TDB or BJD without any scale specified. However, not all of the works cited by Ivshina & Winn (2022) have properly identified their timing systems, nor have they always converted between timing systems accurately. We have identified multiple errors in this work (see Section 6.3 and Section 6.2) relating to missing or erroneously applied BJD/TDB corrections for data that were compiled by Ivshina & Winn (2022).
For all of our new light curves, we begin with data from the instrument in JD/UTC times and use the astropy.time package to convert to BJD/TDB directly. For ETD light curves, we assume that the times listed in the summary table for each planet have been properly converted to HJD/UTC, and apply the correction from UTC to TDB using the same astropy functions. It would be preferable to use the JD times directly from the ETD and convert to BJD, but it is currently not possible to download JD times and observatory coordinates in bulk and is impractical to do so by hand for thousands of light curve times. Thus while most of our times are in BJD/TDB, the ETD times are in a hybrid HJD/TDB system. This has minimal impact given the size of the ETD midtime errors.
@@ 5.4 Calculating the best linear and quadratic ephemerides for each planet
Because of the complexity of the datasets involved, our timing analysis took an iterative approach as we identified errors in the source data. We began with the series of transit midtimes and errors from both our own fits (section 4) and from our curated collection of literature times (section 5). For each system, we then ran the following sequence of calculations, the results of which are illustrated in Figures 3 to 6:
@@
Calculated the value (Equation 5) to see which model is preferred. Negative values of indicate a linear ephemeris is preferred, while positive values indicate a quadratic ephemeris is preferred. The sign of determines if the system had a period increase (positive ) or decrease (negative ).
@@
Ran a series of omit-one tests, omitting each individual midtime in turn and calculating the resulting . Erroneous mid-transit times can significantly impact , particularly if they have very small error bars that act to anchor to a particular model. Individual points whose absence or presence changes by more than 25% were flagged and show up as red diamonds in Figures 3 to 9.
@@
Calculated the rescaled value. Some, though not all, transit midtimes are reported with unrealistically low error bars. The ideal practice would be to refit all literature light curves with the same fit model and method; however, this is usually impractical or impossible, especially for large datasets where the photometry for many light curves have not been published. Instead, we devised a ”rescaling test”, where we used the value of the best linear fit using the original errors and then scaled all error bars up by the same factor, so that , then ran a separate fit to the rescaled data. We then calculated the resulting value using the rescaled error bars. Figure 2 shows the impact of rescaling, which almost always decreases the value for , even causing some values to become negative (switching to a preference for a linear ephemeris); however, genuine detections such as WASP-12 b remain highly significant (rescaled vs original ).
@@
Investigated outliers, especially those that failed the omit-one test. If any times needed to be removed or altered, we then re-ran the above.
@@ 5.5 Identifying systems of interest
| [ms yr] | ||
@@
6 Timing-analysis results for individual systems
@@ 6.1 WASP-12 b: solid detection of decreasing orbital period
Timing data for the planet with the clearest case for orbital decay, WASP-12 b (Yee et al., 2020; Baluev et al., 2019; Patra et al., 2017), with d, are shown in Figure 3. Six new full or partial transits were observed with IoIO for this project in 2022-2023 (Table 1). We also fit two sectors of TESS data (Sector 44 and Sector 45) that were released after the analysis of Ivshina & Winn (2022), which had Sector 20 and Sector 43. We used 145 light curves from the literature as compiled by Ivshina & Winn (2022), plus 28 additional, non-duplicated light curves from the ETD (requiring midtime errors less than five minutes and DQ=1) (Poddaný et al., 2010). The best-fit parameters for linear and quadratic ephemerides are shown in Table 6. Despite the modest precision of the new data (due mostly to unfavorable weather), our value for ms yr is quite similar to previously published values, e.g. ms yr from Yee et al. (2020).
@@ 6.2 WASP-19 b: no evidence for orbital decay
Our analysis includes 2 new transits from the SAAO 40-inch in 2022 as well as two newly-fit sectors of TESS data (Sectors 62 and 63). We used 145 light curves from the literature as compiled by Ivshina & Winn (2022) (but see below for corrections that we made), plus 28 additional, non-duplicated light curves from the ETD, requiring midtime errors less than five minutes and DQ=1 (Poddaný et al., 2010). With the corrected literature times, we find with statistically equal to zero. Decay, if it is happening for WASP-19b, is proceeding at an as-yet undetectable pace.
6.2.1 Errors in the literature: times are not always faithfully propagated
We included WASP-19 b in this paper because using the unmodified literature data as compiled by Ivshina & Winn (2022), it had a highly significant rate of period decrease, with and a 4 detection of decreasing dP/dE. On closer examination, however, we uncovered at least three separate sources of clear-cut errors in the literature, resulting from improper handling of transit midtimes in one of several original works.
@@ 6.2.1 Errors in the literature: times are not always faithfully propagated
The strength of the initial value rested entirely on four points which failed our omit-one test; removing each point individually reduced the value for the system by 30-60%, and removing all of them sent it close to 0. Here is where the chain of custody of values becomes important: we sourced these times from Ivshina & Winn (2022), whose Table 4 reproduced the values from Patra et al. (2020), who used midtimes from transits originally observed and reported by Espinoza et al. (2019). The transits were observed with the 6.5 m Magellan/IMACS, and were of such high precision that even with visible spot crossings during transit the achieved timing precision was better than 10 s. (These low timing errors are why the points had such a large effect on the measured .) However, the midtimes values in published in Patra et al. (2020) do not match the midtimes in Espinoza et al. (2019). The Espinoza et al. (2019) midtimes were published in the BJD/TDB timing system, but differ from the values of Patra et al. (2020) by 67 seconds for the transits taken in 2014-2015, and 69 seconds for the transits taken in 2017. Since that amount is the difference in converting from UTC to TDB, we suspect that an additional, unnecessary correction was applied by Patra et al. (2020), which was then propagated to Ivshina & Winn (2022).
@@
For some reason, Patra et al. (2020) only re-published midtimes for five of the six transits observed by Espinoza et al. (2019), which were discussed above in Item 1. The sixth midtime from Espinoza et al. (2019) was reported as (BJD/TDB). Two transit midtimes at that epoch are reported in Ivshina & Winn (2022) and sourced to Petrucci et al. (2020), who noted that one was from a light curve reported to the ETD and one from the transit observed by Espinoza et al. (2019); it is unclear which midtime was which, since the values reported in Petrucci et al. (2020) do not match the times in either of the original sources. In this case, the origin of the difference in timing (about 40 s in both cases) between the original source and the reanalysis of Petrucci et al. (2020) is unclear but likely represents a separate timing conversion error. We use the original times in Espinoza et al. (2019) and in the ETD in preference to the erroneously compiled values.
3.
A third set of discrepancies was found for three transit midtimes reported originally by Tregloan-Reed et al. (2013) as , 2455252.58506, and 2455255.74045 in the HJD/UTC system. Patra et al. (2020) included these times in their re-analysis and converted them to BJD/TDB, which values are reported in Ivshina & Winn (2022). However, in this case their conversion appears to have not been done correctly. We used an online tool666https://astroutils.astronomy.osu.edu/time/hjd2bjd.html that can convert from HJD (in either UTC or TT) to BJD/TDB with an accuracy of about 1 s using the method in (Eastman et al., 2010). Doing so, we found values for the three converted times (BJD/TDB: 2455251.79706, 2455252.58584, and 2455255.74123) that differ by 35-42 s from the times reported in Patra et al. (2020), and which we use in these analyses. Since these transits were very closely spaced in time, the values from Patra et al. (2020) appear as a single thick red x around E=600 in Figure 4, while our corrected points appears as a thick black dot directly above them.
6.2.2 What if multiple fits to the same light curve yield large differences in errors?
We also found some less clear-cut cases that do not involve errors, but do involve judgement calls over which transit light curve fit (and resulting values and errors) should be used. Consider three precisely timed transits that were observed by Sedaghati et al. (2017). The table from Ivshina & Winn (2022) cites the midtime values for those transits from Petrucci et al. (2020), which uniformly re-fit 74 light curves of WASP-19 b, including these three. The values fit were nearly the same, but the error bars decreased dramatically, e.g., in the original vs. the re-fit – less than 2 sec difference in midtime but a factor of three smaller error. We cannot determine in this work which of the two fits has the more realistic error bars, but we can state that the choice of which midtime error to use has a large effect on the value for . In this case, drops by a third if we choose to use the midtime and larger error from Sedaghati et al. (2017). We have chosen to use the larger error as the more conservative choice.

### v2

@@ 5.2.1 Some examples of composite times in this work
In Section 6.4 we discuss WASP-121 b, where a single composite point, in this case an ephemeris time, is primarily responsible for the nonzero value. In that case, we do not have access to either the individual times or the associated light curves that were used to calculate , making its impact hard to evaluate.
CoRoT-2 b, meanwhile, contains an example of a stacked midtime. The time series from the CoRoT satellite mission encompasses 82 consecutive, high-quality individual transits, but the data were first published as a single stacked midtime assigned to the midtime of the first transit, with [BJD] (Alonso et al., 2008). In this work we reanalyzed the original photometry so that we could individually fit each of the 82 transits (see Table 5), which had an average error of -0.00020, +0.00023 days in the timing of each transit (50% larger than the error assigned to the single stacked midtime). In an experiment, we found significant shifts in the preferred ephemeris model using the stacked midtime rather than the 82 individual midtimes: fell from 33 to and went from 2- positive curvature to negative curvature that is indistinguishable from 0. In this particular case, since each of the individual transits could be well fit to a light curve model on its own, there is no need to use a stacked transit midtime.
We also found that the CoRoT-2 b mission data were being double-counted in Ivshina & Winn (2022). Both the stacked composite time (derived from all 82 transits) from Alonso et al. (2008) and 79 of the 82 individual transit light curves are included in the table in Ivshina & Winn (2022). This duplication had two causes: (1) The individual transit midtimes came from Öztürk & Erdem (2019), who did not analyze the first three transits taken at the longer 512 s cadence; but since the stacked transit was assigned to the midtime of the first (omitted) transit, it appeared to be at a unique epoch. (2) A typo in Ivshina & Winn (2022) caused the omission of the “” tag to indicate that the Alonso et al. (2008) time was derived from more than one transit. In this case, double-counting the CoRoT data had only a modest downward effect on (decreasing it by 6).
@@ 5.3 Some words of caution on timing systems
Eastman et al. (2010) provide a thorough discussion of the various timing systems in use in modern astronomy. Briefly, to remove the effects of both the motion of the Earth around the Sun and the motion of the Sun around the solar system barycenter, a dynamical time system must be used to avoid introducing cyclical variations in an astronomical time series. The recommended standard for reporting exoplanet transit times is Barycentric Julian Dates (BJD) using the Barycentric Dynamical Time (TDB) scale. Critically, for precise timing analyses, all times must be converted into the same timing system before analysis so as to avoid introducing systematic errors.
@@
The difference between the solar system barycenter and the heliocenter is small (no more than a few seconds; Eastman et al., 2010) and BJD and HJD have sometimes been used interchangeably, since almost no transit midtimes are known to better than a few seconds. This practice would reasonable if the times were also both in the same dynamical system – but BJD times are almost always given in the TDB system, while HJD times are usually (though not always) in the UTC system. The difference between UTC and TDB is currently 69.184 s (in 2024). The UTC-TDB offset changes with the irregular addition of leap seconds that depends on the precise rotation of the Earth, and the last leap second was added on 2016 December 31555https://www.nist.gov/pml/time-and-frequency-division/time-realization/leap-seconds. Thus it is critical to report the full time system used, and herein lie some difficulties with literature light curves. Often, times are reported as BJD only, with BJD/TDB implied, though occasionally light curves will be reported in BJD/UTC. The table of transits from Ivshina & Winn (2022) reports the timing system used for each light curve, usually either BJD/TDB or BJD without any scale specified. However, not all of the works cited by Ivshina & Winn (2022) have properly identified their timing systems, nor have they always converted between timing systems accurately. We have identified multiple errors in this work (see Sections 6.3 and 6.2) relating to missing or erroneously applied BJD/TDB corrections for data that were compiled by Ivshina & Winn (2022).
For all of our new light curves, we begin with data from the instrument in JD/UTC times and use the astropy.time package to convert to BJD/TDB directly. For ETD light curves, we assume that the times listed in the summary table for each planet have been properly converted to HJD/UTC and apply the correction from UTC to TDB using the same astropy functions. It would be preferable to use the JD times directly from the ETD and convert to BJD, but it is currently not possible to download JD times and observatory coordinates in bulk, and it is impractical to do so by hand for thousands of light curve times. Thus while most of our times are in BJD/TDB, the ETD times are in a hybrid HJD/TDB system. This has minimal impact given the size of the ETD midtime errors.
@@ 5.4 Calculating the best linear and quadratic ephemerides for each planet
Because of the complexity of the datasets involved, our timing analysis took an iterative approach as we identified errors in the source data. We began with the series of transit midtimes and errors from both our own fits (section 4) and our curated collection of literature times (section 5). For each system, we then ran the following sequence of calculations:
@@
Calculated the value (Equation 5) to see which model is preferred. Negative values of indicate a linear ephemeris is preferred, while positive values indicate a quadratic ephemeris is preferred. The sign of determines if the system had a period increase (positive ) or decrease (negative ). Values of for each planet are shown in Figure 2.
@@
Ran a series of omit-one tests, omitting each individual midtime in turn and calculating the resulting . Erroneous mid-transit times can significantly impact , particularly if they have very small error bars that act to anchor to a particular model. Individual points whose absence or presence changes by more than 25% were flagged and show up as red diamonds in Figures 3 to 10.
@@
Calculated the rescaled value. Some, though not all, transit midtimes are reported with unrealistically low error bars. The ideal practice would be to refit all literature light curves with the same fit model and method; however, this is usually impractical or impossible, especially for large datasets where the photometry for many light curves has not been published. Instead, we devised a ”rescaling test”, where we used the value of the best linear fit using the original errors and then scaled all error bars up by the same factor, so that , then ran a separate fit to the rescaled data. We then calculated the resulting value using the rescaled error bars. Figure 2 shows the impact of rescaling, which almost always decreases the value for , even causing some values to become negative (switching to a preference for a linear ephemeris); however, genuine detections such as WASP-12 b remain highly significant (rescaled versus original ).
@@
Investigated outliers, especially those that failed the omit-one test. If any times needed to be removed or altered, we then reran the above.
@@ 5.5 Identifying Systems of Interest
| [ms yr-1] | ||
| Constant Orbital Periods | ||
| Parameter | CoRoT-2b | WASP-19b |
| Linear P (d) | ||
| Linear | ||
| Quad. P (d)b | ||
| Quad. b | ||
| Quad. dP/dEb | () | ( |
| [ms yr-1] | ||
@@
| b. Best fit quadratic ephemeris for comparison; use linear ephemeris for prediction. | ||
6 Timing Analysis Results for Individual Systems
@@ 6.1 WASP-12 b: solid detection of decreasing orbital period
Timing data for the planet with the clearest case for orbital decay, WASP-12 b (Patra et al., 2017; Baluev et al., 2019; Yee et al., 2020), with d, are shown in Figure 3. Six new full or partial transits were observed with IoIO for this project in 2022-2023 (Table 1). We also fit two sectors of TESS data (Sector 44 and Sector 45) that were released after the analysis of Ivshina & Winn (2022), which had Sector 20 and Sector 43. We used 145 light curves from the literature as compiled by Ivshina & Winn (2022), plus 28 additional, nonduplicated light curves from the ETD (requiring midtime errors less than five minutes and DQ=1, Poddaný et al., 2010). The best-fit parameters for linear and quadratic ephemerides are shown in Table 6. Despite the modest precision of the new data (due mostly to unfavorable weather), our value for ms yr-1 is quite similar to previously published values, e.g. ms yr-1 from Yee et al. (2020).
@@ 6.2 WASP-19 b: no evidence for orbital decay
Our analysis includes 2 new transits from the SAAO 40-inch in 2022 as well as two newly-fit sectors of TESS data (Sectors 62 and 63), for a total of 74 new transit midtimes.
We started with 141 literature times for WASP-19 b listed in Ivshina & Winn (2022). Two points were rejected for having errors greater than five minutes. We then identified 10 times from Cortés-Zuleta et al. (2020) that were not included in Ivshina & Winn (2022), of which we omitted two: one had midtime error greater then five minutes, and the other is a non-obvious duplicate. The duplicate was found in Table A.5 of Cortés-Zuleta et al. (2020), where two nominally different epochs are reported at E=-2063 and from Hebb et al. (2010) and E=-2061 and from Anderson et al. (2010). On closer examination the midtime at E=-2061 is derived from a joint analysis of an occultation observed on 2009 May 3 by Anderson et al. (2010) and the same transit light curve from 2008 Nov 11 from Hebb et al. (2010) that was reported at E=-2063, and thus does not represent an independent midtime measurement. (Occultation timing analysis is beyond the scope of the present work.) We also downloaded 84 midtimes from the ETD (through Feb 2024). Of these, 34 midtimes had DQ=1 and errors less than five minutes (Poddaný et al., 2010). This number was reduced to 29 unique transits after removing two midtimes that were entered from the literature (Hebb et al., 2010; Dragomir et al., 2011), two midtimes from lightcurves that were later reanalyzed and published by Petrucci et al. (2020) and then Ivshina & Winn (2022), and one identical duplicate in the ETD table. Our total number is thus 176 literature midtimes.
We then made three small modifications to the literature record. (1) We restored an extra significant figure to the values from Ivshina & Winn (2022), which were rounded to 5 places, using the original values published in Petrucci et al. (2020). This avoids misidentifying as duplicates a few transits that were observed in multiple filters using the same instrument due to having very similar midtimes but otherwise does not impact results. (2) For one transit originally published by Espinoza et al. (2019), we made a different choice than Ivshina & Winn (2022) about which of several published midtimes to use. All six transits published by Espinoza et al. (2019) also appear with identical values and errors in Patra et al. (2020) after converting to BJD/TDB, but just one appears, non-identically, in Petrucci et al. (2020) where the light curve was reanalyzed; this was the value that appeared in Ivshina & Winn (2022). The difference between the two midtimes in question is 26 seconds, and Petrucci et al. (2020) also found much higher error bars (13 s vs. 4 s). Since only one of the six Espinoza et al. (2019) transits was reanalyzed by Petrucci et al. (2020), and since the Espinoza et al. (2019) times are critical points for the analysis (about which more soon) we made the decision to treat the six transits consistently and used all six BJD/TDB midtimes as published in Patra et al. (2020). (3) For the three times from Sedaghati et al. (2017), we swapped in the original values and errors for the ones that Petrucci et al. (2020) published using re-fit light curves and which are used by Ivshina & Winn (2022). The midtime values were nearly the same, but two of the three light curves have notably smaller error bars in Petrucci et al. (2020), especially the transit in the original vs. the refit – less than 2 sec difference in midtime but a factor of three smaller error. We have chosen to use the larger error as the more conservative choice, although the ultimate effect on after correcting the other issues noted below is minimal.
Using the essentially unmodified literature data as just described we find , which is quite high, with a statistically significant , a nominal 8-sigma detection (see Figure 4). However, closer inspection reveals several issues that bring this result into doubt, which we will now discuss in more detail, along with the impact of each on the calculated value.
@@
First, we found an apparent timing conversion error in Tregloan-Reed et al. (2013), which has since been propagated to other works (including Patra et al., 2020; Ivshina & Winn, 2022). Two tables in Tregloan-Reed et al. (2013) contain timing information, with their Table 2 listing the fitted midtimes for three transit light curves in HJD/UTC and their Table 4 converting to HJD/TDB for direct comparison with other works. At the observation epoch the difference between these two time systems should be about 66 seconds (accounting for both leap seconds and the TAI constant offset), but the actual differences between the listed midtimes are 41 s, 33 s, and 38 s. The cause for this difference is unknown, but may be an incomplete timing conversion, where either the leap seconds or the TAI offset were included, but not both. Using the online tool from Eastman et al. (2010)666https://astroutils.astronomy.osu.edu/time/hjd2bjd.html. Starting from HJD times carries a disclaimer in that tool: “Due to the approximations inherent to most HJD calculators, this conversion cannot be trusted to better than 1 second.” to directly convert the times in their Table 2 from HJD/UTC to BJD/TDB, we find values of 2455251.797060, 2455252.585840, and 2455255.741230, which are used in all subsequent analyses. The uncorrected points are highlighted in Figure 4 as the orange squares around E=600.
Fixing the timing of these three points alone caused a 25% drop to .
@@
After correcting the timing for the three times from Tregloan-Reed et al. (2013), the strength of the value rests on three points which fail our omit-one test; removing each point individually changes the value for the system by 30%. Two of these times are among the six transits discussed above from Espinoza et al. (2019). These transits were white light transits created by combining high-resolution spectroscopic observations with the 6.5 m Magellan telescope using IMACS, and were of such high precision that even with visible spot crossings during transit the timing precision was between 4-9 s in the original work. Besides the low errors, all six points from Espinoza et al. (2019) appear above the line for the best linear ephemeris by about a minute (see the purple triangles in Figure 4). It is worth noting that while at first glance this might appear to be a UTC/TDB conversion error, we have not found any definitive evidence to support that hypothesis. The values in Espinoza et al. (2019) were published in BJD/UTC, as confirmed by correspondence with the authors of that work, and apparent differences between the values in Espinoza et al. (2019) and subsequent works are due to correctly converting the original midtimes to BJD/TDB. The source of the apparent universal offset is thus still unknown.
In combination, the very low errors and the consistent offset are entirely responsible for all of the remaining . We explored two approaches to dealing with this issue:
Increasing the errors on the six points from Espinoza et al. (2019). Noting that the single transit that was re-fit by Petrucci et al. (2020) had three times the fitted error as that reported by Espinoza et al. (2019), we tripled the error bars on all six transits and recalculated the value of . This caused a drop to insignificance, with and . However, one of the Sedaghati et al. (2017) points remains an outlier in this approach.
Removing the six points from Espinoza et al. (2019). This causes the effect to vanish with and , a strong preference for a linear ephemeris. Also, with this approach there are no critical times that affect the value by more than 25%. This is the version we adopted as preferred and which will be used in subsequent analyses in this paper. With these modifications to the literature times, we find that decay, if it is happening for WASP-19 b, is proceeding at an as-yet undetectable pace.